OpenStudy (anonymous):

Please help 2x+6y= 5 x+3y=2 Howmany solutions (x,y) are there to the system of equations above?

OpenStudy (anonymous):

perhaps multiply equation 2 by a factor of 2 and see what you get

OpenStudy (mrnood):

first: multiply the second equation by 2 and post the result here...

OpenStudy (anonymous):

we know whats up haha

OpenStudy (jdoe0001):

what's up? the ceiling, then clouds, then the stratophere

OpenStudy (anonymous):

these are parallel lines

OpenStudy (jdoe0001):

I'd say, solve each equation for "y" so you can see their slope same slopes, no solutions since they're parallel different slopes, one solution

OpenStudy (anonymous):

@MrNood and @jdoe0001 its 2x+6y=4?

OpenStudy (anonymous):

we would clearly see the troposphere before the stratosphere..hahaa

OpenStudy (jdoe0001):

@chris00 and after that, some asteroids floating about, make sure you wear a helmet there though :P

OpenStudy (mrnood):

ok - that's correct so if you look at the first equation 2x + 6y=5 and your equivalent to the seconf 2x+6y=4 then you can see that they cannot BOTH be true for the same values of x and y therefore there is no solution to th e'system' the lines do not intersect

OpenStudy (anonymous):

Okay thanks I remember now @MrNood :)

OpenStudy (anonymous):

parallel and distinct lines